HUBBLE SPACE TELESCOPE DAILY REPORT #1280 PERIOD COVERED: 0000Z Daily Status Report as of 341/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5446 (The Nuclei of Nearby S0 and Spiral Galaxies: A PC Snapshot Imaging) The WF/PC-2 was used to make observations of the core region of galaxy NGC-4600 to provide statistics on the nuclear cups, stellar condensations, and inner stellar disks and dust lanes. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 5560 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1) The WF/PC-2 was used to make several flat calibration observations to monitor the CCD camera electronics, the optical alignment, and to check for contamination. The observations were completed as planned, and no anomalies were noted. 1.3 Completed S/C 5582 (FOC/48 Monthly Activation (Camera Section Only): Cycle 4) This proposal was executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 5393 (The Nature of the Host Galaxies of High-Redshift, Radio-Loud Quasars) The WF/PC-2 was used to make observations of galaxy PKS2338+04 which contains a high red-shifted, radio-loud quasar to study its morphology. This data will be compared with low-Z QSOs and other radio galaxies. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Two Sets of FOS/RD 5535 (Cycle 4 Internal Wavelength Calibration) The Faint Object Spectrograph was used to make internal calibration observations to monitor the stability of the wavelength scale for different gratings. The observations were completed as planned, and no problems were reported. 1.6 Completed GHRS/2 5769 (UV Spectroimaging of Saturn and Search for Exogenous Water in the Atmosphere) The High Resolution Spectrograph ( Side Two) was used to make ultraviolet spectrographic observations of the North and South Hemispheres of the Planet Saturn to look for exogenous water. The target acquisition was performed on the moon Mimus, followed by offsets to several regions on Saturn. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed WF/PC-2 5764 (WF/PC-2 Cycle 4 Calibration: Internal Flats - Repeat) The WF/PC-2 was used to make internal flat calibration images using the VIS lamp and internal sources. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed WF/PC-2 5629 (WF/PC-2 Cycle 4 Photometric Calibration Monitor 1C: 4-Chip STD in F170W) The WF/PC-2 was used to make photometric calibration observations of the ultraviolet standard star GRW+70D582. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2 5479 (Subarcsecond Structure in Nearby AGNs) The WF/PC-2 was used to make snapshot observations of the nearby active galactic nuclei of NGC-6217 to study the interactions between the central engine, the stars, and the interstellar matter in the AGN. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed GHRS/1 5177 (Physical Conditions in Low Z Absorption Line Systems in QSOs: Cycle 4) The High Resolution Spectrograph (Side One) was used to make spectroscopic observations of the QSO PG1630+377 to look for absorption lines in the line of sight of the quasar. Of particular interest is the helium line at 584 A0, which will be shifted to 1443 A0 due to the redshift of Z = 1.473. The observations were executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Significant HSTARs: None 2.4 Operations Notes: Yesterday, HST began the transition to the Second TDRS Ground Terminal (STGT) with the activation of operations with the West Satellite. Most services have been uneventful, but there have been four instances of engineering data loss during two STGT supports. In each case a switch was made to WSGT/NGT, and the problems ceased. These instances were reported to the network via TTRs which are being investigated. Today, the East Satellite will be added to the operational mix. Early this morning (341/1036Z), the Gyro #5 motor encountered a current jump. After the initial spike went to 288 mA, the gyro 5 motor current returned to 156 mA and has remained at that value. In addition to the motor current jump, the gyro bias has changed. All other gyro telemetry looks nominal. There will be a meeting today at 1PM to review actions. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-